# Demystifying Platform Requirements for Diverse LLM Inference Use Cases

### v1

@@ Abstract
Large language models (LLMs) have shown remarkable performance across a wide range of applications, often outperforming human experts. However, deploying these parameter-heavy models efficiently for diverse inference use cases requires carefully designed hardware platforms with ample computing, memory, and network resources. With LLM deployment scenarios and models evolving at breakneck speed, the hardware requirements to meet SLOs remains an open research question. In this work, we present an analytical tool, GenZ, to study the relationship between LLM inference performance and various platform design parameters. Our analysis provides insights into configuring platforms for different LLM workloads and use cases. We quantify the platform requirements to support SOTA LLMs under diverse serving settings. Furthermore, we project the hardware capabilities needed to enable future LLMs potentially exceeding hundreds of trillions of parameters. The trends and insights derived from GenZ can guide AI engineers deploying LLMs as well as computer architects designing next-generation hardware accelerators and platforms. Ultimately, this work sheds light on the platform design considerations for unlocking the full potential of large language models across a spectrum of applications. The source code is available at GitHub.
@@ I Introduction
Large Language Models (LLMs) are among the most popular variants of generative AI models performing auto-regressive tasks. LLMs are used in various applications, including chatbots, code generation, question answering, chemical property prediction[79], protein sequencing[49], and video generation[93, 30]. Commercial products like ChatGPT[54], Gemini[69], Claude[45], Github Copilot[104], have performed astonishingly well in diverse metrics, often outperforming human experts[115]. LLMs have shown great scaling law property [88, 52], with larger models [50, 45, 59] demonstrating better performance as compared to the smaller ones [125]. Currently, the largest model is known to have $\sim$1.8T parameters [29], with future LLMs potentially exceeding even a few hundred trillion parameters.
@@
LLM inference can be broken into two stages, namely the prefill and the decode stage. The prefill stage consists of a single forward pass using all the input tokens. The auto-regressive decode stage generates one output token with each forward pass of the model. The prefill stage often portrays characteristics similar to training forward pass and has significant computing requirements. In contrast, the decode stage requires memory bandwidth and capacity (especially when processing long context queries).
To provide efficient LLM inference, extensive efforts are being made to develop optimizations from both the industry and academic community, including model compression (quantization [68, 86, 129], tensor decomposition [92], pruning [66, 47, 83], knowledge distillation[70, 75]), fast token decoding [91, 67, 64, 57]), system optimizations (paged attention[96], flash attention[61], chunked inference[76], scheduling[130]), and so on. Many of these techniques have also become part of the inference-serving engine such as TensorRT-LLM [105], vLLM [96], DeepSpeed-FastGen [76], and inference servers such as RayLLM [28] with RayServe, TGI[31] and TensorRT-LLM + Triton[39].
Figure 1: Summary of platform requirements for two workloads.
The aforementioned works have enabled vast inference performance improvements for LLMs on current hardware platforms [^1] We define platform as the complete hardware back-end including NPUs (e.g., GPUs or TPUs), local memories (e.g. HBM), remote offload memories (e.g., CPU DRAM) and inter-NPU interconnect (e.g., NVlinks/XeLinks).. Unfortunately, they do not automatically provide an answer to how to design next-generation platforms for LLM inference. It is not trivial as there are a plethora of emerging LLMs with increasing accuracy and context lengths [133], a diverse suite of LLM inference use cases, and various AI platform vendors (e.g., NVIDIA, Intel, AMD, Google, Cerebras, and others) with their own HW/SW frameworks.
The aim of this work is to demystify and quantify AI platform requirements (i.e., compute FLOPs, memory capacity, memory bandwidth, interconnect latency, and interconnect bandwidth) across a suite of LLM inference use cases and representative models. To this end, we introduce a first-order analytical model, namely GenZ  (Generative LLM analyZer), to distill down the impact of different hardware parameters on end-to-end inference performance. Using this tool, we can study current platforms via efficiency factors for the hardware components (which we determine from real-system validation measurements) without getting limited by their specific software tool chains. We also use GenZ to quantify the hardware requirements for different LLM use cases. These studies help us gain first-order vendor-agnostic insights into platform design and configurations.
To the best of our knowledge, this is the first work to quantify platform-wide requirements for meeting the serving SLOs of generative AI models for different use cases. We believe this work makes two unique contributions to the community. First, the data provided by this work can potentially guide AI engineers working on deploying LLMs to choose suitable hardware platform configurations for their use case(s). Next, this work provides concrete directions to computer architects designing the next generation of AI platforms by navigating the interplay between various HW characteristics and LLM inference performance.
Fig. 1 highlights a subset of our key observations for two distinct deployment scenarios. We have quantified and noticed a significant difference in the serving requirements between the prefill and decode stages. Although the absolute values fluctuate with the workload, the identified trend remains consistent. We identify FLOPS and interconnect BW as the bottleneck for the prefill stage, while memory BW and interconnect link latency are the key bottlenecks for the decode stage.
@@
To summarize, this work makes the following contributions:
  • We introduce GenZ, a first-order analytical tool that helps analyze LLM workload on different platforms. Fig. 2 shows an overview of our proposed tool.
  • We provide an analysis of the performance impact of different workload optimizations using GenZ.
  • We provide an analysis of the performance impact of different platform characteristics using GenZ.
  • We provide an analysis of the platform requirements for the target LLM inference to meet the given Service Level Objectives (SLOs) with the target use case using GenZ.
II Key Takeaways
Following are the key takeaways from our work. For detailed understanding please read the rest of the paper. You can also use our code to regenerate the results.
  • The prefill and decode stages of the LLM inference process have significant differences in their workloads.
  • We find that increasing the TFLOPS and inter-chip bandwidth reduces the prefill latency but doesn’t impact the decoding latency.
  • On the other hand, increasing the high-speed memory bandwidth(HBM) and reducing the inter-chip network link latency helps improve the latency of the decode stage but does not reduce the latency of the prefill stage.
  • We also find that for larger systems parallelism plays a major role, where in the prefill stage, using pipeline parallelism gives the best throughput. During the decode stage, a mix of tensor and pipeline parallelism is optimal for throughput in large systems and tensor parallel for smaller systems. .
  • We also quantify the requirements for meeting the SLO throughput requirements of various models with various use cases and provide analytical formulas to modify the demand according to the usecase and model changes.
@@ III-A LLM Architecture
Mixture of Experts (MoE)[65] are special class of LLMs that consist of multiple “expert” multi-layer perceptron (MLP) layers, denoted as ‘E’, out of which ‘K’ experts are selected for each input token. In contrast, dense language models can be considered as a special case of MoEs, where E = K = 1, meaning there is only one expert MLP layer, and it is used for every input token. By introducing multiple experts and selectively activating a subset of them for each token, MoEs can effectively scale model capacity while maintaining efficient computation and memory usage, making them a promising approach for building larger and more capable language models. Some popular MoE models are Switch Transformers[65], Mixtral 8x7B[85], Mixtral 8x22B[22], DBRX[5], GROK[11], GPT-4[41, 9, 10]. Table I presents some of the state-of-the-art LLMs and associated dimensions.
@@
LLaMA2-7B[123] 4096 32 32 $\sim$2.7 -
Mixtral-8x7B[85] 4096 32 32(8) 3.5 8(2)
@@
Super LLM-10T 13824 128 108 4 32(4)
TABLE I: Different LLM architecture evaluated in this work.
III-B Generative LLM Inference
@@
Prefill. The prefill stage runs only once on input sequence of $\tau_{p}$ tokens to generate the K and V activations, which are often kept as the KV cache, for each LLM layer. The generated KV cache would be used for all subsequent output token generation. Table II shows the operator wise breakdown of each LLM decode layer during the prefill stage. $C_{op}$ and $M_{op}$ is the number of floating point calculations, memory requirement required for the operator. The prefill stage is mostly compute-bounded as all input tokens can be processed in parallel.
Operation Input Shapes Output Shape $C_{op}$ $M_{op}$
$Q/K/V$ [B, $\tau_{p}$, $\underline{D_{model}}]\times$ [$\underline{D_{model}},3D_{model}$] 3B, $\tau_{p},H,d$ $O(3B\tau_{p}D_{model}^{2})$ $O(4B\tau_{p}D_{model}+3D_{model}^{2})$
$Q*K^{T}$ [B, $\tau_{p},H,\underline{d}]\times[B,\tau_{p},H,\underline{d}]$ B, H, $\tau_{p}$, $\tau_{p}$ $O(BH\tau_{p}^{2}d)$ $O(2B\tau_{p}D_{model}+BH\tau_{p}^{2})$
$P*V$ [B, $\tau_{p}$, $\underline{\tau_{p}}]\times[B,\underline{\tau_{p}}$, H, d] B, H, $\tau_{p}$, d $O(BH\tau_{p}^{2}d)$ $O(2B\tau_{p}D_{model}+BH\tau_{p}^{2})$
$A*W_{O}$ [B, $\tau_{p}$, $\underline{D_{model}}]\times$ [$\underline{D_{model}},D_{model}$] B, $\tau_{p}$, $D_{model}$ $O(B\tau_{p}D_{model}^{2})$ $O(B\tau_{p}D_{model}+D_{model}^{2})$
$FF_{up}$ [B, $\tau_{p}$, $\underline{D_{model}}]\times$ [$\underline{D_{model}},D_{ff}$] B, $\tau_{p}$, $D_{ff}$ $O(B\tau_{p}D_{model}D_{ff})$ $O(B\tau_{p}D_{model}+D_{model}D_{ff}+B\tau_{p}D_{ff})$
$FF_{gate}$ [B, $\tau_{p}$, $\underline{D_{model}}]\times$ [$\underline{D_{model}},D_{ff}$] B, $\tau_{p}$, $D_{ff}$ $O(B\tau_{p}D_{model}D_{ff})$ $O(B\tau_{p}D_{model}+D_{model}D_{ff}+B\tau_{p}D_{ff})$
$FF_{down}$ [B, $\tau_{p}$, $\underline{D_{ff}}]\times$ [$\underline{D_{ff}},D_{model}$] B, $\tau_{p}$, $D_{model}$ $O(B\tau_{p}D_{model}D_{ff})$ $O(B\tau_{p}D_{ff}+D_{ff}D_{model}+B\tau_{p}D_{model})$
TABLE II: Matrix Multiplication operations during the prefill stage. Reduction dimensions are underscored. Blue colored operators are model weights.
@@
Decode. After the prefill stage, output tokens are generated auto-regressively, i.e. the last generated token is fed as an input to the LLM in each iteration, and one new output ‘token’ is generated. Table III shows the operator-wise breakdown of each LLM decode layer during the decode stage. $C_{op}$ and $M_{op}$ is the number of floating point calculations, memory requirement required for the operator. All the model weights and past KV cache are used to generate a single output token. Since input to the model is a single token, all matrix-matrix multiplications are converted into matrix-vector multiplications. This makes the generation stage highly memory-bounded. At the end of each generation step, the newly generated KV cache is padded to the existing KV cache, meaning its linear growth with the output sequence length. Generation stops when a special $<$end-token$>$ is generated, or the user limits the maximum number of output tokens. We define $\tau_{d}$ tokens are generated during the decode phase.
Operation Input Shapes Output Shape $C_{op}$ $M_{op}$
$Q/K/V$ [B, 1, $D_{model}$] $\times$ [$\underline{D_{model}},3D_{model}$] B, H, 1, d $O(3BD_{model}^{2})$ $O(3D_{model}^{2})$
$Q*K^{T}$ [B, $1,H,\underline{d}]\times$ $[B,\tau_{cur}+1,H,\underline{d}]$ B, H, 1, $\tau_{cur}+1$ $O(BH(\tau_{cur}+1)d)$ $O(B\tau_{cur}D_{model})$
$P*V$ [B, 1, $\underline{\tau_{cur}+1}]\times$ [B, $\tau_{cur}+1$, H, d] B, H, 1, d $O(BH(\tau_{cur}+1)d)$ $O(B\tau_{cur}D_{model})$
$A*W_{O}$ [B, 1, $\underline{D_{model}}]\times$ [$\underline{D_{model}},D_{model}$] B, 1, $D_{model}$ $O(BD_{model})$ $O(D_{model}^{2})$
$FF_{up}$ [B, 1, $\underline{D_{model}}]\times$ [$\underline{D_{model}},D_{ff}$] B, 1, $D_{ff}$ $O(BD_{ff}D_{model})$ $O(D_{ff}D_{model})$
$FF_{gate}$ [B, 1, $\underline{D_{model}}]\times$ [$\underline{D_{model}},D_{ff}$] B, 1, $D_{ff}$ $O(BD_{ff}D_{model})$ $O(D_{ff}D_{model})$
$FF_{down}$ [B,1, $\underline{D_{ff}}]\times$ [$\underline{D_{ff}},D_{model}$] B, 1, $D_{model}$ $O(BD_{ff}D_{model})$ $O(D_{ff}D_{model})$
TABLE III: Matrix Multiplication operations during the decode stage. Reduction dimensions are underscored. Blue and red colored operators are model weights and KV caches respectively. $\tau_{cur}$ is sum of prefill tokens($\tau_{p}$) and token generated by LLM before the current token.
Beam search [121, 89, 77, 103] is a heuristic search algorithm commonly employed in LLMs during the decode stage by recent LLMs. It helps generate more contextually relevant and coherent sequences of words by exploring multiple potential sequences simultaneously. The number of maximum parallel sequences is a user-defined parameter called Beam Size, $S_{b}$. Typically $S_{b}=4$ for generative inference. During generative LLM inference, the model predicts the next word based on previous context with beam search maintaining a limited set of the likely candidate sequences. These candidate sequences are ranked by their likelihood, and the algorithm continually refines them by considering alternative word choices. By exploring a predefined number of possibilities in parallel, beam search enhances the quality of the generated text. After the common prefill stage, two parallel beams are launched during the decode phase. The beam with the highest likelihood is selected as the final output.
III-C Retrieval Augmented Generation
Despite the promise of being an exciting paradigm towards Artificial General Intelligence (AGI), LLMs still suffer from the obstacles of factual hallucination [53, 112, 84, 71], incomplete, outdated or incorrect knowledge [73], and the lack of domain-specific knowledge [100, 118]. Enhancing LLMs with external knowledge through Retrieval Augmented Generation (RAG) has been proven effective against all such challenges [97, 72, 48, 80] as mentioned in [56].
Before prefill, RAG takes an input query embedding, compares it against a large vector database using KNN search, and augments the query with the k most similar vectors retrieved from the database. As the size of these knowledge databases can be of the order of trillions [48], several ANN (approximate nearest neighbor) techniques including IVF-PQ[82], HNSW[102], SONG[134], and CAGRA[106] have been proposed over the years that trade-off search latency and memory footprint with recall accuracy. The key implication of using RAG in LLM serving is a significant increase in the input prompt size. This results in higher Time Per Output Token (TPOT) and larger KV caches. We study these implications in detail in Section VII.
III-D Metrics for LLM Serving

### v2

# Demystifying AI Platform Design for Distributed Inference of Next-Generation LLM models

@@ Abstract
Large language models (LLMs) have shown remarkable performance across a wide range of applications, often outperforming human experts. However, deploying these gigantic models efficiently for diverse inference use cases requires carefully designed hardware platforms with ample computing, memory, and network resources. With constant innovation in LLM serving optimizations and model architecture evolving at breakneck speed, the hardware requirements to meet Service Level Objectives (SLOs) remain an open research question. To answer the question, we present an analytical tool, GenZ, to efficiently navigate the relationship between diverse LLM model architectures, LLM serving optimizations, and AI platform design parameters. Our tool estimates LLM inference performance metrics for the given scenario. We have validated against real hardware platforms running various different LLM models, achieving a max geomean error of 5.82%. We use GenZ to identify compute, memory capacity, memory bandwidth, network latency, and network bandwidth requirements across diverse LLM inference use cases. We also study diverse architectural choices in use today (inspired by LLM serving platforms from several vendors) to help inform computer architects designing next-generation AI hardware accelerators and platforms. The source code is available at GitHub. GenZ can also be tried out on its website without any setup on your web browser.
@@ I Introduction
The success of LLMs has fueled a growing interest in Generative AI use cases - spanning Question-Answer bots, text summarization, code generation, image generation, video generation, and more. Commercial products like ChatGPT, Gemini, Github Copilot  [1, 2, 3], have performed astonishingly well in diverse metrics, often outperforming human experts  [4]. LLMs have shown great scaling law properties [5, 6], with larger models  [7, 8, 9] demonstrating better performance as compared to smaller ones  [10]. Currently, the largest model has $\sim$1.8T parameters  [11], and future LLMs could potentially have even a few hundred trillion parameters.
@@
The design of LLM serving systems has become a hot area of research. This is due to its unique computational characteristics that set it apart from traditional Deep Learning inference and training. LLM serving involves two distinct stages: prefill and decode. The prefill stage consists of a single forward pass using all the input tokens. This is followed by an auto-regressive decode stage that generates one output token with each forward pass of the model. The prefill stage often portrays characteristics similar to a traditional forward pass (inference) and has significant computing requirements. In contrast, the decode stage consumes (and generates) one token at a time and leverages a large cache of Key and Value projections of the input tokens, requiring high memory bandwidth (BW) and capacity (especially when processing long context queries). The metrics for LLM serving are also unique - with the use case playing a key role in determining the latency criticality of prefill vs decode tokens.
Figure 1: Platform requirements for two workloads.
@@
Framework LLM Architecture Supported Model Level Optimizations* System Level Optimizations Distributed Platform Modeling
Parallelism Network Hierarchy Memory Hierarchy
ASTRA-sim  [12] Dense None TP, PP Multi-Level
Multi-Level (HBM
+ DRAM offload)
LLM-Viewer  [13] Dense, Dense-GQA Flash-attention, quantization* TP Single Level Single-Level (HBM)
LLM-Compass  [14] Dense None TP, PP Single Level Single-Level (HBM)
Vidur  [15] Dense, Dense-GQA Flash-attention, chunked prefill TP, PP Real HW only Single-Level (HBM)
LLMServingSim  [16] Dense Flash-attention, quantization* TP, PP Multi-Level Single-Level (HBM)
GenZ (This Work)
Dense, Dense-GQA
MoE, Mamba
Sliding window attention
Flash-attention, chunked prefill, beam-search
speculative decoding, quantization*, sparsity*
mixed precision computation*, KV pruning*
TP, PP, EP Multi-Level
Multi-Level (HBM
+ DRAM offload)
TABLE I: Comparison of prior works for modeling LLM inference against this work. TP, PP, and EP identify as Tensor, Pipeline, and Expert parallelism, respectively.
To state that LLM inference is an active area would be an understatement. The use cases for LLM inference continue to grow by the day. This is fueled by two trends. First, new LLM models with enhanced accuracy are being released at a rapid cadence by multiple competing AI labs  [2, 17, 18, 19, 20, 21]. Second, for each model, a plethora of model-level (i.e., lossy) optimizations (such as quantization  [22, 23, 24], pruning  [25, 26, 27], fast token decoding  [28, 29, 30, 31])) and system-level (i.e., lossless) optimizations (such as paged attention  [32], flash attention  [33], chunked inference  [34, 35], scheduling  [36]) are employed for enhancing performance by reducing the compute and memory footprint. Many of these optimization techniques have also become part of the popular inference-serving engine such as TensorRT-LLM  [37], vLLM  [32], and DeepSpeed-FastGen  [34].
While GPU-based scaled-out platforms [^1] We define platform as the complete hardware back-end including multiple NPUs (e.g., GPUs or TPUs), local memories (e.g. HBM) and inter-NPU interconnect (e.g., NVlinks/XeLinks). have gone mainstream for large model training, the jury is out on the right architectural platform for LLM inference. Today, there exist several architecturally distinct platforms that have all shown competitive performance across multiple LLM use cases [38]. Examples include GPU-based platforms from NVIDIA and AMD, programmable scaled-out dataflow processors from SambaNova, and SRAM-only architectures from Cerebras (via waferscale) and Groq (via hundreds of interconnected ASICs). There is also a lot of excitement around a plethora of emerging LLM platforms - such as transformer-specific ASICs from Etched [39] and high-speed photonics from Lightmatter [40].
We aim to demystify and quantify design insights for AI platforms across a suite of LLM model architectures, system optimizations and use cases. To this end, we introduce GenZ (Generative LLM analyZer), a framework for modeling and evaluating the relationship between LLM uses cases, model optimizations, software optimizations and the hardware platform and predict the end-to-end LLM inference performance. It should be noted that GenZ does not intend to simulate individual NPUs, rather we focus on simulating the distributed platforms for various LLM architectures combined with inference optimizations. While there exist valuable tools in the community to design NPUs [41, 42, 43] and distributed network fabric architectures [12], GenZ is the first tool, to the best of our knowledge, capturing the full-spectrum of LLM inference optimizations enabling the isolation and study of specific model/software/hardware optimizations on the end-to-end LLM performance (or energy). Table I contrasts GenZ against related efforts on distributed platform modeling. In fact, GenZ is able to plug-in external tools for compute and communication to enable high-fidelity modeling of the underlying hardware. We validate GenZ thoroughly with various LLMs served across different platforms. Promisingly, across different workloads, GenZ can closely mimic the performance on diverse real platforms, achieving a geomean error of only 5.82%.
We demonstrate the value of GenZ via multiple case studies, answering several open questions surrounding hardware requirements stemming from diverse model types, software optimizations, and architectural choices across diverse use-cases. For instance, Fig. 1 highlights a subset of our key observations for two distinct scenarios.
In summary, this work makes the following contributions:
1
We introduce GenZ, an analytical framework that helps analyze LLMs combined with various inference optimizations on different platforms (Section III). Fig. 2 shows an overview of our proposed framework. 2 Using GenZ, we study the impact of various LLM serving optimizations with current hardware specifications, and present key insights to guide the next-generation design specifications (Section IV). 3 We present detailed runtime analysis across various LLM architectures based on both the transformer and state-space models (Section V). 4 Lastly, we showcase GenZ’s modeling capability through four case studies on the next-generation AI inference platform design: (1) isolated scaling of different platform characteristics (Section VII-A); (2) comparison of different AI platform architecture design choices (for example, SRAM-Chiplet, Wafer-Scale, GPUs, ASICs) (Section VII-B); (3) study design space choices of high bandwidth domain (HBD) size as well as associated interconnects (Section VII-C); and (4) study the impact of micro-architecture details and compute offloading(Section VII-D).
@@ II-A LLM Architecture
Mixture of Experts (MoE)[45] are special class of LLMs that consist of multiple “expert” multi-layer perceptron (MLP) layers, denoted as ‘E’, out of which ‘K’ experts are selected for each input token. In contrast, dense language models can be considered as a special case of MoEs, where E = K = 1, meaning there is only one expert MLP layer, and it is used for every input token. By introducing multiple experts and selectively activating a subset of them for each token, MoEs can effectively scale model capacity while maintaining efficient computation and memory usage, making them a promising approach for building larger and more capable language models. Some popular MoE models are Switch Transformers[45], Mixtral 8x7B[46], Mixtral 8x22B[47], DBRX[48], GROK[20], GPT-4[49, 50, 17]. Table IV presents some of the state-of-the-art LLMs and associated dimensions.
@@
Gemma2-2B [51] 2304 26 8(4) 4 -
LLaMA3-8B [52] 4096 32 32(8) 3.5 -
Gemma2-27B [51] 4608 46 32(16) 8 -
Mixtral-8x22B [47] 6144 56 48(8) 2.66 8(2)
@@
Dense-5T 49152 128 192(24) 4 -
MoE-10T 13824 128 108(12) 4 32(4)
TABLE II: Different LLM architectures evaluated in this work.
II-B Generative LLM Inference
@@
Prefill. The prefill stage runs only once on input sequence of $\tau_{p}$ tokens to generate the K and V activations, which are often kept as the KV cache, for each LLM layer. The generated KV cache would be used for all subsequent output token generation. The prefill stage is mostly compute-bounded as all input tokens can be processed in parallel.
@@
Decode. After the prefill stage, output tokens are generated auto-regressively, i.e. the last generated token is fed as an input to the LLM in each iteration, and one new output ‘token’ is generated. All the model weights and past KV cache are used to generate a single output token. Since input to the model is a single token, all matrix-matrix multiplications are converted into matrix-vector multiplications. This makes the generation stage highly memory-bounded. At the end of each generation step, the newly generated KV cache is padded to the existing KV cache, meaning its linear growth with the output sequence length. Generation stops when a special $<$end-token$>$ is generated or the user limits the maximum number of output tokens. We define $\tau_{d}$ tokens are generated during the decode phase.
Chunked Prefill  [35, 53, 54, 34] or chunking is a recent serving optimization used to reduce the hardware under-utilization by combining the two stages of LLM generation to provide better throughput. In chunked prefill, a chunk of fixed size is fed to the model in each iteration. All existing decode batches are processed parallelly in the chunk. The remaining slots are filled by outstanding prefill requests. Since the prefill is broken down into smaller chunks combined with the decode stage, this ameliorates the memory boundedness of the decode stage and improves the overall throughput of the system. This often comes at slightly increased latency for the prefill stage.
Beam search  [55, 56] is a heuristic search algorithm used during the decoding stage. It explores multiple potential sequences simultaneously, with the number of parallel sequences, called the beam size ($S_{b}$), typically set to 4. After the common prefill stage, two parallel beams are launched during the decode phase. The beam with the highest likelihood is selected as the final output.
II-C Metrics for LLM Serving